OpenStudy (anonymous):

Which of the following is not a function: A) 2y^2 - 5 B) y = 2sqrt(x) C) y = 2x^2 - 5 I think it is A but i am not sure. Can anyone tell me if it is?

OpenStudy (babyslapmafro):

A

OpenStudy (babyslapmafro):

A is not a relation of two inputs, i.e. (y and x)

OpenStudy (anonymous):

Thank you!

OpenStudy (babyslapmafro):

no prob

OpenStudy (anonymous):

one more thing. would (-1,-6) (4,-6)(-3,-6) be a function? From what I understand if there are any x values that are the same it cant possibly be a function..but what about y?

OpenStudy (babyslapmafro):

I don't understand that question, can you phrase it exactly how it is in the book or online program your are getting the question from?

OpenStudy (anonymous):

Which of the following is a function: A) (-1,-6) (4,-6)(-3,-6) B) x = y^2 C) X^2 + y^2 = 1

OpenStudy (babyslapmafro):

Tricky question, A could represent a function since there are no x-values given with a similar y-value. Also, B is a function since x=y^2 is the same as y=sqrt(x). And C could be a function given similar rules as seen in answer B.

OpenStudy (babyslapmafro):

I would really have to focus on the strict definition of a function to help you answer this multiple choice question, what I have given is all I know at this time, so I won't be able to help much.

OpenStudy (anonymous):

alright well thanks!